OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation. 3x - 5y + z = 12 x + 6y - z = -4 2x - 5y - 2z = -8 (4 points) Answer 4x + y = 8 -17y = 0 2x - 11y = 16 4x + 7y = -16 -23y + 4z = 24 -17y = 0 13y - 2z = 24 7y - 4z = 0

OpenStudy (anonymous):

First, lets solve the second equation for x. What do you get for that?

OpenStudy (anonymous):

what do you mean slove for x i honestly dont have any idea how to do this problem :/

OpenStudy (anonymous):

I'm going to walk you through it. Solve for x means get x by itself on one side of the equal sign. For example, if I were to solve x + 10 + 2x = 0 for x, I would get 3x = -10 then x = -10/3. That being said, the second equation is x + 6y - z = -4. Solve it for x.

OpenStudy (anonymous):

good teaching eashmore.